I have 5 homework problems I am really struggling with. Since 5 is the limit per day, I'll post them all here:
Solve each inequality and graph its solution.
1) 6x - 2x > -12
2) 5k + 4k ≥ 0
3) 6 – 2n +3 > -1
4) 3n – n > 12
5) –(1 + 6x) ≤ 2 – 3x


Answer by Alan3354(69443)

Solve each inequality and graph its solution.
1) 6x - 2x > -12
4x > -12
Divide by 4
---------------------------
2) 5k + 4k >= 0
Combine the 2 terms.
Divide by the coefficient.
-------------------------
3) 6 – 2n +3 > -1
You couldn't add the 6 & 3?
9 - 2n > -1
Subtract 9
-2n > -10
Divide by 2
-n > -5
Add 5
-n + 5 > 0
Add n
5 > n
n < 5
------------------
4) 3n – n > 12
See above.
---------------------
5) –(1 + 6x) <= 2 – 3x
-1 - 6x <= 2 - 3x
See above.